Subject: Raith will remain off line -- likely through Tuesday Noon. UPDATE Monday afternoon .
From: James Conway <jwc@snf.stanford.edu>
Date: Fri, 27 Feb 2004 17:25:41 -0800
Fri, 27 Feb 2004 17:25:41 -0800
Hello Raith Users:

The Raith system will remain off line through next Tuesday Noon at the 
earliest -- please adjust your schedules accordingly.

The Raith computer can be used for pattern editing and preparing 
position list.  Please be careful the when you login that the beam is 
kept unblanked, as the normal default for most user logins is to blank 
on login. If you encounter difficulties call me immediately.

NO ACCESS OR COMMUNICATIONS TO THE LEO COLUMN.  PLEASE LET THE SYSTEM 
IDLE ALONG AND PLEASE DO NOT LOGOUT OF THE SERVICE WINDOWS AND GUN 
MONITORS AS WE ARE COLLECTING  DATA.

Currently we are idling at 20 keV, BEAM ON, EHT ON, 10000 kX mag. Beam I 
= 0.212 nA 

Beam emission has stabilized nicely but the vacuum levels and imaging 
resolution are very poor.
Several sections of the LEO electron optics board were damaged when the 
system arced during user write early Wed. evening.  This was not the 
fault of any user, just the luck of the draw...

LEO and RAITH have provided excellent support and we all worked hard to 
efficiently coordinate and ship out E/O boards and  parts needed to 
effect repairs.  Best efforts were made by all parties involved.  This 
situation will improve early next week.

Thank you for your support,

James Conway

I will be traveling and 'off the cell grid' intermittently through the 
weekend.  415-412-4825











Hello Raith Users:

The Raith system will remain off line through next Tuesday Noon at the earliest -- please adjust your schedules accordingly.

The Raith computer can be used for pattern editing and preparing position list.  Please be careful the when you login that the beam is kept unblanked, as the normal default for most user logins is to blank on login. If you encounter difficulties call me immediately.

NO ACCESS OR COMMUNICATIONS TO THE LEO COLUMN.  PLEASE LET THE SYSTEM IDLE ALONG AND PLEASE DO NOT LOGOUT OF THE SERVICE WINDOWS AND GUN MONITORS AS WE ARE COLLECTING  DATA.

Currently we are idling at 20 keV, BEAM ON, EHT ON, 10000 kX mag. Beam I = 0.212 nA 

Beam emission has stabilized nicely but the vacuum levels and imaging resolution are very poor.
Several sections of the LEO electron optics board were damaged when the system arced during user write early Wed. evening.  This was not the fault of any user, just the luck of the draw...

LEO and RAITH have provided excellent support and we all worked hard to efficiently coordinate and ship out E/O boards and  parts needed to effect repairs.  Best efforts were made by all parties involved.  This situation will improve early next week.

Thank you for your support,

James Conway

I will be traveling and 'off the cell grid' intermittently through the weekend.  415-412-4825